If Olivia Garcia is single and in the 25% tax bracket, calculate the tax associated with each of the following transactions. (Use the IRS regulations for capital gains in effect in 2014.)

EXHIBIT 3.2  Capital Gains Tax Categories as of 2014
Capital gains tax rates are as low as 0 percent for low-income or as high as 28 percent for higher income levels and certain types of assets, so long as the holding period is more than 12 months.
Holding period Tax Brackets/Assets Sold (2014) Tax on Capital gains
Less than 12 months All (10%, 15%, 25%, 28%, 33%, 35%, and Same as ordinary income
39.6%) - any asset sold
Over 12 months 10% and 15% - assets other than real estate and collectibles 0%
25%, 33%, 35%, 39.6% 15%, or 20% (later only if in 39.6% tax bracket)
Sale of depreciable real estate 25% on gain up to depreciation amount
Collectibles 28% all gains from collectibles

Treat each of the following cases as independent of the others.

  1. She sold stock for $3,900 that she purchased for $3,000 6 months earlier. Round the answer to the nearest cent. Tax savings should be preceded by a "-" sign.
    $   
  2. She sold bonds for $4,400 that she purchased for $3,500 3 years earlier. Round the answer to the nearest dollar. Tax savings should be preceded by a "-" sign.
    $  
  3. She sold stock for $2,940 that she purchased for $3,500 20 months earlier. Assume this to be the only Stock in Olivia's portfolio. Round the answer to the nearest cent. Tax savings should be preceded by a "-" sign.
    $  
Answer #1

a). Profit = selling price - purchase price = 3,900-3,000 = 900

Sold within 12 months so short-term capital gains tax will apply. She is within the 25% tax bracket so tax will be 25%*900 = 225.00

b). Profit = selling price - purchase price = 4,400-3,500 = 900

Sold after 12 months so long-term capital gains tax will apply. For a person in the 25% tax bracket, the applicable capital gains tax is 15% so tax to be paid = 15%*900 = 135.00

c). Loss = selling price - purchase price = 2,940-3,500 = -560

Long-term capital tax applicable = 15%*-560 = -84.00 (tax saving)